OpenStudy (anonymous):

Newton's Law of Cooling A Cup of coffee placed in a room cools from 160F to 150F in one minute. If the room temperature is 65F, Find the temperature in ten minutes.

OpenStudy (anonymous):

"Newton's Law of Cooling"\[ T(t) = (T_0-T_r)e^{kt} + T_r \]Correct me if I'm wrong on this formula. "A Cup of coffee placed in a room cools from 160F to 150F in one minute."\[ T(0) = T_0 = 160^\circ F \\ T(1) = 150^\circ F \](This will help us solve for \(k\)) "the room temperature is 65F"\[ T_r = 60 ^\circ F \]"Find the temperature in ten minutes"\[ T(10) = ? \]

OpenStudy (anonymous):

@zhengcl86 Think you can take it from here?

OpenStudy (anonymous):

@wio i have a different formula

OpenStudy (anonymous):

Tr=60∘F why is it 60? and not 65?

OpenStudy (anonymous):

That is typo... what is your formula?

OpenStudy (anonymous):

Xk - Xk-1 = A (Xk-1 - 65)

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

so i am using somewhat a similar formula then yours but i still get an answer

OpenStudy (anonymous):

so i am using a different formula that i found online T(t) = T0 + ( T1 - T0) e^-k(t) T(t) = 65 +(160-65)e^-k(t) T(t) = 65+95 e^-k(t) time = 1 Temp = 150 150 = 65+95 e^-k(1) 85 = 95 e^-k .8947368421 = -k T(t) 65+95 e^-.8947368421(10) T(10) = 160 - 8.947368421 T = 151 which still makes no sense!! what am i doing wrong?

OpenStudy (anonymous):

don't know what i did wrong but the temperature went up 1 degrees in 10 minutes lol

OpenStudy (anonymous):

My formula was perfectly fine.

OpenStudy (anonymous):

\[\begin{array}{} T(t) &=& (T_0-T_r)e^{kt} + T_r \\ 150 &=& (160-65)e^{k(1)}+65 \\ 85 &=& 95e^{k} \\ \frac{17}{19} &=& e^{k} \\ \end{array} \]Plugging it back in: \[\begin{array}{} T(t) &=& 95\left(\frac{17}{19}\right)^{t}+65 \\ T(10) &=& 95\left(\frac{17}{19}\right)^{10}+65 \\ &\approx & 96 \end{array} \]

OpenStudy (anonymous):

how is the work done in T(10)

OpenStudy (anonymous):

I plugged it into wolfram

OpenStudy (anonymous):

oh.....well thats not gonna help me during the test lol thanks you...

OpenStudy (anonymous):

You're not allowed to use a calculator?

OpenStudy (anonymous):

Just do the order of operations.

OpenStudy (anonymous):

i am, im trying to do the calculations now and its not giving me 96

OpenStudy (anonymous):

so 17/19 x 10 = 8.947368421

OpenStudy (anonymous):

that multple by 95 is 850

OpenStudy (anonymous):

Not multiplication. It's an exponent.

OpenStudy (anonymous):

then 850 + 65 is 915

OpenStudy (anonymous):

i am still very lost on the last part

OpenStudy (anonymous):

Exponents are repeated multipicaton... \[ x^{10} = \underbrace{x\times x \times x \times ... \times x}_{10\text{ times}} \]

OpenStudy (anonymous):

where does T(10) go?

OpenStudy (anonymous):

so with the exponent of 17/19 to the 10 is 8.005540166

OpenStudy (anonymous):

\[ \left(\frac{17}{19}\right)^{10}\approx 0.329 \] http://www.wolframalpha.com/input/?i=%5Cleft%28%5Cfrac%7B17%7D%7B19%7D%5Cright%29%5E%7B10%7D

OpenStudy (anonymous):

dont know why i am not getting that on my calculator

OpenStudy (anonymous):

well thank you for all your help